OpenStudy (anonymous):

Part 1: Show the expanded form and simplification of x^4/x^2. Explain in your own words how you can simplify x^4/x^2 without having to write the expanded form. Part 2: Create your own fraction with like bases, coefficients, and show its simplification.

OpenStudy (anonymous):

 ^^^^^ how it

OpenStudy (anonymous):

looks!

OpenStudy (anonymous):

\[\frac{x^4}{x^2}=x^{4-2}=x^2\]